OpenStudy (anonymous):

A history teacher gives a 17-question true-false exam. In how many ways can the test be answered if the possible answers are T or F or possibly to leave the answer blank? 129,140,163 ways is the answer, but I don't know how to get to the answer.

jimthompson5910 (jim_thompson5910):

You compute 2^17 to get that result

jimthompson5910 (jim_thompson5910):

oh wait, sry, 3^17 because you have T, F or B for blank

jimthompson5910 (jim_thompson5910):

the reason being because you have 3 choices for question 1, 3 for question 2, etc etc multiply 3's 17 times to get 3*3*3*....*3*3 = 3^17 = 129,140,163

OpenStudy (anonymous):

That makes sense! Thank you very much!

jimthompson5910 (jim_thompson5910):

you're welcome